3x-y=4 find the slope?
3x-1/5y=-4 Find the slop
2/3y=-7/4x find the slope

Answer by Alan3354(69443)

3x-y=4 find the slope?
3x-1/5y=-4 Find the slop
2/3y=-7/4x find the slope
--------------
2/3y=-7/4x find the slope
----
(2/3)y = (-7/4)x
Solve for y
2y = (-21/4)x
y = (-21/8)x
The slope is the coefficient of x, after solving for y.
m = -21/8
================
Do the others the same way.